# Diffusion differentiable resampling

### v2

@@ Abstract
This paper is concerned with differentiable resampling in the context of sequential Monte Carlo (e.g., particle filtering). We propose a new informative resampling method that is instantly differentiable, based on an ensemble score diffusion model. We theoretically prove that our diffusion resampling method provides a consistent resampling distribution, and we show empirically that it outperforms the state-of-the-art differentiable resampling methods on multiple filtering and parameter estimation benchmarks. Finally, we show that it achieves competitive end-to-end performance when used in learning a complex dynamics-decoder model with high-dimensional image observations.
Machine Learning, ICML, diffusion model, particle filter, sequential Monte Carlo, resampling, differentiable, Feynman–Kac
@@ 1 Introduction
Resampling is a key component in sequential Monte Carlo (SMC) samplers for state-space models (SSMs, Chopin and Papaspiliopoulos, 2020). It not only mitigates particle degeneracy in practice, but has also been shown to represent a jump Markov process in a continuous-time limit (Chopin et al., 2022). However, resampling typically hinders (gradient-based) parameter estimation of SSMs due to discrete randomness. For instance, with the commonly used multinomial resampling, one draws indices $I_{i}\sim\mathrm{Categorical}(w_{1},w_{2},\ldots,w_{N})$ independently for $i=1,2,\ldots,N$, and then define the resampling mapping by indexing $X^{*}_{i}\coloneqq X_{I_{i}}$. If the sample $X^{\theta}_{i}$ (and weight) depend on some unknown parameters $\theta$, the pathwise derivative of the re-sample, $\partial X_{i}^{\theta,*}/\,\partial\theta$, is not defined. Moreover, most automatic differentiation libraries (e.g., JAX) will typically drop the undefined derivatives, resulting in erroneous gradient estimates (see, e.g., Naesseth et al., 2018).
@@
Various methods have thus been proposed to make the resampling step differentiable in the context of SMC and SSMs (see recent surveys in Chen and Li, 2025; Brady et al., 2025). One line of work focuses on the expectation derivative $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$, which is often well defined, rather than the pathwise derivative. This can be achieved by combing Fisher’s score (Poyiadjis et al., 2011) and stochastic derivatives (Arya et al., 2022), resulting in, for example, a stop-gradient based method by Ścibior and Wood (2021). However, these methods often suffer from high variance and may consequently require a large sample size $N$.
@@
Another line of work focuses on developing new resampling methods that naturally come with well-defined pathwise derivatives, in contrast to multinomial and its variants (e.g., residual resampling). Notable examples include soft (Karkus et al., 2018) and Gumbel-Softmax resampling (Jang et al., 2017). Both methods essentially form an interpolation between multinomial resampling (which has no gradient) and uninformative resampling (which has a gradient) via a calibration parameter. Although they have been empirically shown to work well for certain models, they are fundamentally biased. Crucially, one has to decide on a trade-off between the gradient bias and the statistical performance of the resampling mapping. Similarly, Zhu et al. (2020) propose to parametrise the resampling with a neural network, which adds both training complexity and additional sources of bias.
@@
The perhaps first fully-differentiable-by-construction resampling method with consistency guarantees is due to Malik and Pitt (2011), who make a smooth approximation of the empirical cumulative distribution function (CDF) of the weighted samples. However, their method is applicable to univariate $X$ only. This was later generalised by Li et al. (2024) using kernel jittering to approximate the CDF gradient. In a similar vein, Corenflos et al. (2021) propose an optimal transport (OT) based resampling method. The idea is to learn a (linear) transportation map between the target distribution $\pi$ and the proposal, and approximate the resampling by an ensemble transformation (Reich, 2013)
@@
where $P^{\varepsilon}\in\mathbb{R}^{N\times N}$ denotes the $\varepsilon$-regularised entropic optimal coupling. The main problem of OT-based resampling lies in computation, that one has to compute for the transportation plan $P^{\varepsilon}$. The cost scales quadratically in the number of samples $N$ with aSinkhorn implementation, which in turn depends exponentially on the entropy parameter $1\,/\,\varepsilon$ (Luo et al., 2023a; Burns and Liang, 2025) to converge. In the context of SMC, the method may not work well when the proposal/reference does not well approximate the target. Li et al. (2024, Fig. 1) also show a (hypothetical) case when linear transformation of OT is insufficient for exploiting the distribution manifold. Nevertheless, this line of work has inspired us to develop a new transportation-based resampling method that avoids these issues.
Other than making the resampling differentiable, one can also modify the SMC algorithm itself to produce smooth estimates of the marginal log likelihood of SSMs. For instance, Klaas et al. (2005) introduce a mixture of SMC proposals to marginalise out the need for resampling, however, in practice, drawing samples from mixture distributions typically requires discrete sampling too. This was addressed by Lai et al. (2022) using implicit reparametrisation, but the approach remains restricted to structured proposals. Overall, this class of methods rely on customised SMC samplers, potentially limiting their applicability in general.
@@
  • We introduce diffusion resampling, a new resampling paradigm that instantly enables automatic differentiation for $\partial X_{i}^{\theta,*}/\,\partial\theta$, and consequently also for the expectation $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$. We apply the method for filtering and gradient-based parameter estimation in state-space models with SMC samplers.
@@ 3 Weight $w_{0,i}\leftarrow G_{0}^{\theta}(Z_{0,i})\,/\,L_{0}(\theta)$
6    $\bigl\{Z_{j-1,i}^{*}\bigr\}_{i=1}^{N}\leftarrow\texttt{diffres}\Bigl(\bigl\{(w_{j-1,i},Z_{j-1,i})\bigr\}_{i=1}^{N}\Bigr)$
@@ 2.2 Mean-reverting Gaussian reference
To combine with the SMC in Algorithm 2, we compute $\mu_{N}$ and $\Sigma_{N}$ based on $\bigl\{(w_{j-1,i},Z_{j-1,i})\bigr\}_{i=1}^{N}$, see Appendix A for details. It is also possible to leverage any Gaussian filter which is commonly used for approximating the optimal proposal (van der Merwe et al., 2000), to establish the reference. The central gist here is to exploit the sequential structure of SMC to adaptively and informatively choose the reference, instead of assuming a static and uninformative one, such as $\mathrm{N}(0,I_{d})$. Similar ideas of using mean-reverting SDEs for informative generative sampling have been used in domain-specific applications, such as image restoration (Luo et al., 2023b, 2025).
@@ 3 Convergence analysis
A natural follow-up question is whether we can improve the error bound by eliminating the error due to $p_{T}\approx\pi_{\mathrm{ref}}$, that is, $\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})$ in Proposition 1. This may be achieved by using the technique in Corenflos et al. (2025); Zhao et al. (2025), where a forward-backward Gibbs chain $(p_{T|0},q_{T|0})$ is constructed to avoid explicit sampling from the marginal $p_{T}$. This essentially transforms the error $p_{T}\approx\pi_{\mathrm{ref}}$ into a statistical correlation of the chain. However, since Markov chains typically converge geometrically as well, the required number of chain steps is likely comparable to the diffusion time $T$ (Meyn and Tweedie, 2009; Andrieu et al., 2024). Furthermore, in practice we only have access to the approximate conditional $\widetilde{q}_{T|0}$. Nonetheless, we may still benefit computationally from the fact that a fixed $T$ can allow for moderate discretisation. We leave this for future investigation.
@@ 4.1 Gaussian mixture importance resampling
Consider a target model $\pi(x)\propto\phi(x)\,p(y{\;|\;}x)$ with Gaussian mixture prior $\phi(x)=\sum_{i=1}^{c}\omega_{i}\,\mathrm{N}(x;m_{i},v_{i})$ and likelihood $p(y{\;|\;}x)=\mathrm{N}(y;H\,x,\Xi)$. The posterior distribution $\pi$ is also a Gaussian mixture available in closed form (see details in Appendix F). We use the prior as the proposal to generate importance samples, and then apply the resampling methods to obtain resampled particles. These are compared to samples directly drawn from $\pi$, using the sliced 1-Wasserstein distance (SWD) and resampling variance for the mean estimator. We run experiments 100 times independently with sample size $N=10,000$.
@@
Table 1 shows that the diffusion resampling at $K=128$ gives the best SWD, followed by the baseline multinomial resampling and OT ($\varepsilon=0.3$). We also observe that the diffusion resampling performance highly depends on the choice of integrator and discretisation, which at $K=8$ is not superior than OT or multinomial. In terms of resampling variance for estimating the posterior mean, diffusion resampling is not as performant as OT but is still better than the baseline multinomial. OT seems to have stable resampling variance robust in $\varepsilon$. Gumbel-Softmax and the soft resampling methods are not comparable to diffusion or OT, except when they reduce to multinomial resampling numerically. Further results are given in Appendix F.
@@
The time costs of diffusion resampling and OT are shown in Figure 1, and we have two important observations. The left figure shows that both methods roughly scale polynomially in the sample size $N$, where diffusion $(K=4)$ is better than OT $(\varepsilon=0.8)$. On the right side of the figure, both methods scale linearly in their parameters $K$ and $1\,/\,\varepsilon$, which aligns with the summary in Table 17, and the cross point shows that they have a similar computation at $K\approx 6\,/\,\varepsilon$ when $N=8,192$. We also observe a trend that when $N$ increases, the cross point moves left, showing that diffusion scales better than OT in $N$. Details are given in Appendix G.
@@ 4.2 Linear Gaussian SSM
with parameters $\theta_{1}=0.5$ and $\theta_{2}=1$. For each experiment, we generate a measurement sequence with 128 time steps and apply a Kalman filter to compute the true filtering solution and log-likelihood $L(\theta_{1},\theta_{2})$ evaluated at Cartesian $\Theta=[\theta_{1}-0.1,\theta_{1}+0.1]\times[\theta_{2}-0.1,\theta_{2}+0.1]$. Three errors are evaluated: 1) the error of log-likelihood function estimate $\lVert L-\widehat{L}\rVert^{2}_{2}\coloneqq\int_{\Theta}(L(\theta_{1},\theta_{2})-\widehat{L}(\theta_{1},\theta_{2}))^{2}\mathop{}\!\mathrm{d}\theta_{1}\mathop{}\!\mathrm{d}\theta_{2}$, where $\widehat{L}$ stands for the estimate by a particle filter with resampling; 2) the KL divergence between the particle samples and the true filtering solution; 3) the estimated parameters $\widehat{\theta}$ by L-BFGS compared to the truth under Euclidean norm $\lVert\theta-\widehat{\theta}\rVert_{2}$. All results are averaged over 100 independent runs, with 32 particles used. For details see Appendix H.
@@
Table 2: The errors of loss function, filtering (scaled by $10^{-1}$), and parameter estimation (scaled by $10^{-1}$) for Section 4.2. Divergent NaN results are explained in Appendix H.
@@
We also observe that almost all the resampling methods encounter divergent parameter estimations to some extent (see more statistics in Appendix H). In particular, the divergence of Soft and Gumbel-Softmax is too significant to give a meaningful result, and these entries are thus marked as NaN. The reason is due to the quasi-Newton optimiser L-BFGS-B which heavily depends on stable and accurate gradient estimates (Xie et al., 2020). This shows that the diffusion and OT approaches are more applicable for generic optimisers, such as second-order ones. Therefore, in the next sections we will focus on comparison using first-order gradient-based methods.
@@ 4.3 Prey-predator model
Consider a more complicated Lokta–Voltera model
@@
where the parameters and the non-linear rate function $\lambda$ are detailed in Appendix I. We generated data at time $t\in[0,3]$ discretised by Milstein’s method with 256 steps. We model the dynamics entirely by a neural network, without assuming known the dynamic structure. After the neural network has been learnt, we make 100 predictions from the learnt model and then compare them to a reference trajectory generated by the true dynamics. All results are averaged over 20 independent runs, with $N=64$ particles used.
Figure 3: Root mean square errors of predictions based on learnt prey-predator models over independent runs (scatter points). The best results go to the first two columns from diffusion resampling.
@@
Results are summarised in Figures 3 and 4. The first figure shows that the trained model using the diffusion resampling achieves the lowest prediction error and is significantly more stable than the baselines. Furthermore, as detailed in Appendix I, even diffusion resampling with the worst setting gives around 2.67 prediction error which still outperforms the average best of all other methods. The second figure aligns with the results from the first figure, showing that the training loss with diffusion resampling constantly lower bounds that of the other methods, and is more stable.
@@ 4.4 Vision-based pendulum dynamics tracking
where $Z_{j}\coloneqq\begin{bmatrix}Z^{(1)}_{j}&Z^{(2)}_{j}\end{bmatrix}$ represents the state defined by the angle and angular velocity, $\Delta_{j}$ is the discretisation step, and $\zeta_{j}\sim\mathcal{N}(0,\Delta_{j}\,\Lambda_{\zeta})$ is the process noise. The true observation function $r$ generates $32\times 32$ greyscale images representing the pendulum state. We generate a trajectory of observations $Y_{0:J}$ over steps $j=0,1,\ldots,256$. Crucially, in contrast to the previous example in Section 4.3, we here assume that both the underlying dynamics and the observation model are unknown, and we parametrise both the state transition, $Z_{j}=f_{\theta}(Z_{j-1},\zeta_{j})$, and the decoder, $r\approx r_{\phi}$, by neural networks. Parameters $\theta$ and $\phi$ are learnt by minimising the negative log-likelihood estimated by the particle filter with resampling. For completeness, we run experiments in two regimes: the first uses adaptive resampling under a tempered observation likelihood, and the second is a more challenging setting with resampling at nearly every filtering step. The learnt models are evaluated using the average structural similarity index (SSIM) and peak signal-to-noise ratio (PSNR) on their predicted image sequences.
Our results demonstrate that diffusion resampling integrates effectively into this complex, high-dimensional SMC learning pipeline, enabling stable optimisation and achieving competitive performance relative to state-of-the-art differentiable resampling baselines. Figure 6 shows an accurate and high-fidelity mean image sequence reconstructed from a latent pendulum dynamics model and decoder learnt with diffusion resampling embedded in the optimisation process. The comparative results in Figure 5 show that, among the strongest configurations of each resampling class, diffusion resampling achieves mean SSIM and PSNR at least comparable to the baselines. See Appendix J for further details and results.
@@
Figure 6: Qualitative comparison of the learnt pendulum dynamics (trained end-to-end in the first experiment setting) against the ground truth. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom).

### v3

@@ Abstract
This paper is concerned with differentiable resampling in the context of sequential Monte Carlo (e.g., particle filtering). Drawing on reparametrisation, we propose a new resampling method that is informative and instantly differentiable, based on a training-free diffusion model surrogate. We theoretically prove that our diffusion resampling method provides a consistent resampling distribution, and we show empirically that it outperforms the state-of-the-art differentiable resampling methods on multiple filtering and parameter estimation benchmarks. Finally, we show that it achieves competitive end-to-end performance when used in learning a complex dynamics-decoder model with high-dimensional image observations.
Machine Learning, diffusion models, particle filters, sequential Monte Carlo, resampling, differentiable, reparametrisation, Feynman–Kac
@@ 1 Introduction
Resampling is a key component in sequential Monte Carlo (SMC) samplers for state-space models (SSMs, Chopin and Papaspiliopoulos, 2020). It not only mitigates particle degeneracy in practice, but has also been shown to represent a jump Markov process in a continuous-time limit (Chopin et al., 2022). However, resampling typically hinders (gradient-based) parameter estimation of SSMs due to discrete randomness. For instance, with the commonly used multinomial resampling, one draws indices $I_{i}\sim\mathrm{Categorical}(w_{1},w_{2},\ldots,w_{N})$ independently for $i=1,2,\ldots,N$, and then defines the resampling mapping by indexing $X^{*}_{i}\coloneqq X_{I_{i}}$. If the sample $X^{\theta}_{i}$ (and weight) depend on some unknown parameters $\theta$, the pathwise derivative of the re-sample, $\partial X_{i}^{\theta,*}/\,\partial\theta$, is not defined. Moreover, most automatic differentiation libraries (e.g., JAX) will typically drop the undefined derivatives, resulting in erroneous gradient estimates (see, e.g., Naesseth et al., 2018).
@@
Various methods have thus been proposed to make the resampling step differentiable in the context of SMC and SSMs (see recent surveys in Chen and Li, 2025; Brady et al., 2025). One line of work focuses on the expectation derivative $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$, which is often well defined, rather than the pathwise derivative. This can be achieved by combing Fisher’s score (Poyiadjis et al., 2011) and stochastic derivatives (Arya et al., 2022), resulting in, for example, the stop-gradient based method by Ścibior and Wood (2021). However, these REINFORCE-based methods often suffer from high variance and may consequently require a large sample size $N$.
@@
Another line of work focuses on developing new resampling methods that naturally come with well-defined pathwise derivatives (i.e., reparapemtrisation). Notable examples include soft (Karkus et al., 2018) and Gumbel-Softmax resampling (Jang et al., 2017). Both methods essentially form an interpolation between multinomial resampling (which has no gradient) and uninformative resampling (which has a gradient) via a calibration parameter. Although they have been empirically shown to work well for certain models, they are fundamentally biased. Crucially, one has to decide on a trade-off between the gradient bias and the statistical performance of the forward resampling mapping. In addition, Zhu et al. (2020) propose to parametrise the resampling with a neural network, which adds both training complexity and additional sources of bias. Kviman et al. (2024) introduce a deterministic resampling, which also comes with irreducible biases (Finke et al., 2026).
@@
The perhaps first fully-differentiable-by-construction reparametrisation with consistency guarantees is due to Malik and Pitt (2011), who make a smooth approximation of the empirical cumulative distribution function (CDF) of the weighted samples. However, their method focuses on univariate $X$. This was later generalised by Li et al. (2024) using kernel jittering to approximate the CDF gradient. In a similar vein, Corenflos et al. (2021) propose an optimal transport (OT) based resampling method. The idea is to learn a (linear) transportation map between the target distribution $\pi$ and the proposal, and approximate the resampling by an ensemble transformation (Reich, 2013)
@@
where $P^{\varepsilon}\in\mathbb{R}^{N\times N}$ denotes the $\varepsilon$-regularised entropic optimal coupling. The main problem of OT-based resampling lies in computation, since one has to compute for the transportation plan $P^{\varepsilon}$. The cost scales quadratically in the number of samples $N$ with a Sinkhorn implementation, which in turn depends exponentially on the entropy parameter $1\,/\,\varepsilon$ (Luo et al., 2023a; Burns and Liang, 2025) to converge. In the context of SMC, the method may not work well when the proposal/reference does not well approximate the target. Li et al. (2024, Fig. 1) also show a (hypothetical) case when linear transformation of OT is insufficient for exploiting the distribution manifold. Nevertheless, this line of work has inspired us to develop a new transportation-based resampling method that avoids these issues.
Other than making the resampling differentiable, one can also modify the SMC algorithm itself to produce smooth estimates of the marginal log likelihood of SSMs. For instance, Klaas et al. (2005) introduce a mixture of SMC proposals to marginalise out the need for resampling, however, in practice, drawing samples from mixture distributions typically also requires discrete sampling. This was addressed by Lai et al. (2022) using implicit reparametrisation, but the approach remains restricted to structured proposals. Overall, this class of methods rely on customised SMC samplers, potentially limiting their applicability in general.
@@
  • We introduce diffusion resampling, a new reparametrisation paradigm that instantly enables automatic differentiation for $\partial X_{i}^{\theta,*}/\,\partial\theta$, and consequently also for the expectation $\partial\operatorname{\mathbb{E}}\bigl[X_{i}^{\theta,*}\bigr]/\,\partial\theta$. We apply the method for filtering and gradient-based parameter estimation in state-space models with SMC samplers.
@@ 3 Weight $w_{0,i}\leftarrow G_{0}^{\theta}(Z_{0,i})\,/\,L_{0}(\theta)$
6    $Z_{j-1,i}^{*}\leftarrow\texttt{diffres}\Bigl(\bigl\{(w_{j-1,i},Z_{j-1,i})\bigr\}_{i=1}^{N}\Bigr)$
@@ 2.2 Mean-reverting Gaussian reference
To combine with the SMC in Algorithm 2, we compute $\mu_{N}$ and $\Sigma_{N}$ based on $\bigl\{(w_{j-1,i},Z_{j-1,i})\bigr\}_{i=1}^{N}$, see Appendix A for details. It is also possible to leverage any Gaussian filter which is commonly used for approximating the optimal proposal (van der Merwe et al., 2000), to establish the reference. The gist here is to exploit the sequential structure of SMC to adaptively and informatively choose the reference, instead of assuming a static and uninformative one, such as $\mathrm{N}(0,I_{d})$. Using mean-reverting SDEs for informative generative sampling have also been used in domain applications, such as image restoration (Luo et al., 2023b, 2025).
@@ 3 Convergence analysis
A natural follow-up question is whether we can improve the error bound by eliminating the error due to $p_{T}\approx\pi_{\mathrm{ref}}$. This may be achieved by using the technique in Corenflos et al. (2025); Zhao et al. (2025), where a forward-backward Gibbs chain $(p_{T|0},q_{T|0})$ is constructed to avoid explicit sampling from the marginal $p_{T}$. This essentially transforms the error $p_{T}\approx\pi_{\mathrm{ref}}$ into a statistical correlation of the chain. However, since Markov chains typically converge geometrically as well, the required number of chain steps is likely comparable to the diffusion time $T$ (Meyn and Tweedie, 2009; Andrieu et al., 2024). Furthermore, in practice we only have access to the approximate conditional $\widetilde{q}_{T|0}$. Nonetheless, we may still benefit computationally from the fact that a fixed $T$ can allow for moderate discretisation.
@@ 4.1 Gaussian mixture importance resampling
Consider a target distribution $\pi(x)\propto\phi(x)\,p(y{\;|\;}x)$ with Gaussian mixture prior $\phi(x)=\sum_{i=1}^{c}\omega_{i}\,\mathrm{N}(x;m_{i},v_{i})$ and likelihood $p(y{\;|\;}x)=\mathrm{N}(y;H\,x,\Xi)$. The posterior distribution $\pi$ is also a Gaussian mixture available in closed form (see Appendix G). We use the prior as the proposal to generate importance samples, and then apply the resampling methods to obtain resampled particles. These are compared to samples directly drawn from $\pi$, using the sliced 1-Wasserstein distance (SWD) and resampling variance for the mean estimator. We run experiments 100 times independently with sample size $N=10,000$.
@@
Table 1 shows that the diffusion resampling at $K=128$ gives the best SWD, followed by the baseline multinomial resampling and OT ($\varepsilon=0.3$). We also observe that the diffusion resampling performance highly depends on the choice of integrator and discretisation, which at $K=8$ is not superior than OT or multinomial. In terms of resampling variance for estimating the posterior mean, diffusion resampling is not as performant as OT but is still better than the baseline multinomial. OT seems to have stable resampling variance robust in $\varepsilon$. Gumbel-Softmax and the soft resampling methods are not comparable to diffusion or OT. Further results are given in Appendix G.
@@
The time costs of diffusion resampling and OT are shown in Figure 1, and we have two important observations. The left figure shows that both methods roughly scale polynomially in the sample size $N$, where diffusion $(K=4)$ is better than OT $(\varepsilon=0.8)$. On the right side of the figure, both methods scale linearly in their parameters $K$ and $1\,/\,\varepsilon$, which aligns with the summary in Table 20, and the cross point shows that they have a similar computation at $K\approx 6\,/\,\varepsilon$ when $N=8,192$. We also observe a trend that when $N$ increases, the cross point moves left, showing that diffusion scales better than OT in $N$. Details are given in Appendix H.
@@ 4.2 Linear Gaussian SSM
with parameters $\theta_{1}=0.5$ and $\theta_{2}=1$. For each experiment, we generate a measurement sequence with 128 time steps and apply a Kalman filter to compute the true filtering solution and log-likelihood $L(\theta_{1},\theta_{2})$ evaluated at Cartesian $\Theta=[\theta_{1}-0.1,\theta_{1}+0.1]\times[\theta_{2}-0.1,\theta_{2}+0.1]$. Three errors are evaluated: 1) the error of log-likelihood function estimate $\lVert L-\widehat{L}\rVert^{2}_{2}\coloneqq\int_{\Theta}(L(\theta_{1},\theta_{2})-\widehat{L}(\theta_{1},\theta_{2}))^{2}\mathop{}\!\mathrm{d}\theta_{1}\mathop{}\!\mathrm{d}\theta_{2}$, where $\widehat{L}$ stands for the estimate by a particle filter with resampling; 2) the KL divergence between the particle samples and the true filtering solution; 3) the estimated parameters $\widehat{\theta}$ by L-BFGS compared to the truth under Euclidean norm $\lVert\theta-\widehat{\theta}\rVert_{2}$. All results are averaged over 100 independent runs, with 32 particles used. For details see Appendix I.
@@
Table 2: The errors of loss function, filtering (scaled by $10^{-1}$), and parameter estimation (scaled by $10^{-1}$) for Section 4.2. Divergent NaN results are explained in Appendix I.
@@
We also observe that almost all the resampling methods encounter divergent parameter estimations to some extent (see more statistics in Appendix I). In particular, the divergence of Soft and Gumbel-Softmax is too significant to give a meaningful result, and these entries are thus marked as NaN. The reason is due to the quasi-Newton optimiser L-BFGS-B which heavily depends on stable and accurate gradient estimates (Xie et al., 2020). This shows that the diffusion and OT approaches are more applicable for generic optimisers, such as second-order ones. Therefore, in the next sections we will focus on comparison using first-order gradient-based methods.
@@ 4.3 Prey-predator model
Consider a more challenging Lokta–Voltera model
@@
where the configuration is detailed in Appendix J. We generate data at time $t\in[0,3]$ discretised by Milstein’s method with 256 steps. We model the dynamics entirely by a neural network, without assuming known the dynamic structure. After the neural network has been learnt, we make 100 predictions from the learnt model and then compare them to a reference trajectory generated by the true dynamics. All results are averaged over 20 independent runs, with $N=64$ particles used. We have also compared to the REINFORCE framework implemented with a stopped gradient method by Ścibior and Wood (2021).
Figure 3: Root mean square errors of the learnt prey-predator models over independent runs (scatter points).
@@
Results are summarised in Figures 3 and 4. The first figure shows that the trained model using the diffusion resampling achieves the lowest prediction error and is significantly more stable than the baselines. The second figure aligns with the results from the first figure, showing that the training loss with diffusion resampling consistently lower bounds that of the other methods, and is more stable.
@@ 4.4 Vision-based pendulum dynamics tracking
where $Z_{j}\coloneqq\begin{bmatrix}Z^{(1)}_{j}&Z^{(2)}_{j}\end{bmatrix}$ represents the state defined by the angle and angular velocity, $\Delta_{j}$ is the discretisation step, and $\zeta_{j}\sim\mathcal{N}(0,\Delta_{j}\,\Lambda_{\zeta})$ is the process noise. The true observation function $r$ generates $32\times 32$ snapshots of the pendulum. We generate a trajectory of observations $Y_{0:J}$ over $J=256$ steps. In contrast to Section 4.3, we learn both the underlying dynamics and the observation model, and we parametrise both the state transition, $Z_{j}=f_{\theta}(Z_{j-1},\zeta_{j})$, and the decoder, $r\approx r_{\phi}$, by neural networks. Parameters $\theta$ and $\phi$ are learnt by minimising the negative log-likelihood estimated by the particle filter with resampling. For completeness, we run experiments in two regimes: the first uses adaptive resampling under a tempered observation likelihood, and the second is a more challenging setting with resampling at nearly every filtering step. The learnt models are evaluated using the average structural similarity index (SSIM) and peak signal-to-noise ratio (PSNR) on their predicted image sequences.
Our results demonstrate that diffusion resampling integrates effectively into this complex, high-dimensional SMC learning pipeline, enabling stable optimisation and achieving competitive performance relative to state-of-the-art differentiable resampling baselines. Figure 6 shows an accurate and high-fidelity mean image sequence reconstructed from a latent pendulum dynamics model and decoder learnt with diffusion resampling embedded in the optimisation process. The comparative results in Figure 5 show that, among the strongest configurations of each resampling class, diffusion resampling achieves mean SSIM and PSNR at least comparable to the baselines. See Appendix K for further details and results. We also include additional large-scale experiments in Appendix L and Appendix M.
@@
Figure 6: Qualitative comparison of the learnt pendulum dynamics. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom).